OpenStudy (anonymous):

compute limit x->3 x^4-81/2x^2-5x-3

OpenStudy (lgbasallote):

\[\frac{x^4 - 81}{2x^2 - 5x - 3} \implies \frac{(x^2-9)(x^2+9)}{(2x+1)(x-3)} \implies \frac{(x+3)(x-3)(x^2 + 9)}{(2x+1)(x-3)}\] does that help?

OpenStudy (lgbasallote):

i have to go sleep now so i hope you got that ^_^

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

welcome